### v1

@@ 2 Preliminaries
This section reviews the first-order necessary optimality conditions for MPECs, some tailored stationarity concepts, and some regularization methods for solving MPECs. Both are necessary to state and justify the steps of the algorithm proposed in this paper and to analyze its convergence properties.
2.1 Optimality conditions for MPECs
We start by defining some useful index sets for the complementarity constraint of the MPEC (1). Denote the feasible set of the MPEC (1) as $\Omega=\{x\in{\mathbb{R}}^{n}\mid c(x)\geq 0,0\leq x_{1}\perp x_{2}\geq 0\}$. The following index sets that depend on a feasible point $x$ that satisfy complementarity constraints (1c) are considered:
${\mathcal{I}}_{+0}(x)$ =$=\{i\in\{1,\ldots,m\}\mid x_{1,i}>0,x_{2,i}=0\},$
${\mathcal{I}}_{0+}(x)$ =$=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}>0\},$
${\mathcal{I}}_{00}(x)$ =$=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}=0\}.$
As we will see below, most of the computational and theoretical difficulties arise if the so-called biactive set ${\mathcal{I}}_{00}(x)$ is nonempty. The active set of the standard inequality constraints (1b) is defined as ${\mathcal{A}}(x)=\{i\mid c_{i}(x)=0\}$.
Next, we define first-order necessary optimality conditions. To sate the first-order necessary optimality conditions for a generic NLP, including an MPEC, we need the notion of the tangent cone. The tangent cone at $x\in\Omega$ to the set $\Omega$ is defined as $\mathcal{T}_{\Omega}(x)=\{d\in{\mathbb{R}}^{n}\mid\exists\{x^{k}\}\subset\Omega,\{t^{k}\}\subset{\mathbb{R}}_{\geq 0}:\lim\limits_{k\to\infty}t^{k}=0,\lim\limits_{k\to\infty}x^{k}=x,\lim\limits_{k\to\infty}\frac{x^{k}-x}{t^{k}}=d\}$. Let $d=(d_{0},d_{1},d_{2})\in{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{m}$. The first-order necessary optimality conditions for (1) are given by the following theorem (Luo1996, , Corollary 3.3.1).
@@ 2.1 Optimality conditions for MPECs
Let $x^{*}\in\Omega$ be a local minimizer of (1), then it holds that
@@
or equivalently, $d=0$ is a local minimizer of the following optimization problem:
@@
Note that Theorem 2.1 is purely geometric and thus difficult to verify. In standard optimization theory, to obtain an algebraic characterization of a stationary point, the tangent cone ${\mathcal{T}}_{\Omega}(x)$ is replaced by the linearized feasible cone (Nocedal2006, , Def. 12.3 ). However, the latter is always a convex polyhedral cone, and if ${\mathcal{I}}_{00}(x)$ is nonempty the cone ${\mathcal{T}}_{\Omega}(x)$ is always nonconvex Luo1996 , making this transition impossible. A more suitable notion is the the so-called MPEC-linearized feasible cone Flegel2005a ; Pang1999 ; Scheel2000 , which for a feasible point $x\in\Omega$ reads as
@@
Observe that at a feasible point $x\in\Omega$ all constraint residuals above are zero but they are kept for convenience. In contrast to the classical linearized feasible cone, here the combinatorial structure is kept for the degenerate index set ${\mathcal{I}}_{00}(x)$ and $\mathcal{F}_{\Omega}(x)$ is nonconvex, if $\mathcal{I}_{00}$ is nonempty. It is a union of a finite number of convex polyhedral cones.
It general it holds that ${\mathcal{T}}_{\Omega}(x)\subseteq\mathcal{F}_{\Omega}(x)$ Flegel2005a . But if the so-called MPEC-Abadie constraint qualification is satisfied, it follows that ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$ Flegel2005a and we can use the MPEC-linearized feasible cone in Theorem 2.1. In particular, one can define a linear program with complementary constraints (LPEC) to characterize a (algebraic) B-stationary point.
Definition 2.1 (B-stationariry)
@@
s.t. $c(x^{*})+\nabla c(x^{*})^{\top}d\geq 0,$ (6b)
$x_{1,i}^{*}+d_{1,i}=0,$ $\forall i\in{\mathcal{I}}_{0+}(x),$ (6c)
$x_{2,i}^{*}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x),$ (6d)
$0\leq x_{1,i}^{*}+d_{1,i}\perp x_{2,i}^{*}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{00}(x).$ (6e)
Note that if $d=0$ is a local minimizer of (6), then it follows from the LPEC that $x^{*}\in\Omega$. In summary, to verify that $x^{*}$ is B-stationary, one has to verify that $d=0$ is a local minimizer of an LPEC (or LP), which may be computationally expensive if $|{\mathcal{I}}_{00}(x^{*})|\gg 1$.
There exists also multiplier-based stationarity concepts for MPECs that are defined via the KKT conditions applied to more regular NLPs derived from the MPEC (1). Such problems will also be solved in every iteration of the method proposed in this paper. Thus, we defined them explicitly in the sequel.
Let $x^{*}\in\Omega$ and ${\mathcal{D}}_{1}(x^{*}),{\mathcal{D}}_{2}(x^{*})$ be a partition of the biactive set ${\mathcal{I}}_{00}(x^{*})$, with $\mathcal{D}_{1}(x^{*})\cup\mathcal{D}_{2}(x^{*})={\mathcal{I}}_{00}(x^{*})$ and $\mathcal{D}_{1}(x^{*})\cap\mathcal{D}_{2}(x^{*})=\emptyset$. Define the index sets:
@@
The number of possible partitions at a feasible point $x^{*}$ is equal to $2^{|{\mathcal{I}}_{00}(x^{*})|}$. For every partition, we can define a branch NLP, denoted by $\mathrm{BNLP}{({\mathcal{I}}_{1},{\mathcal{I}}_{2})}$, as follows:
@@
$x_{1,i}\geq 0,\ x_{2,i}=0,\;$ $\forall i\in{\mathcal{I}}_{2}(x^{*}),$ (8d)
If the standard constraints $c(x)\geq 0$ do not lead to a violation of constraint qualifications, (8) are regular NLPs for which the standard optimization theory is applicable and existing numerical algorithms perform well.
Additionally, the MPEC literature frequently uses the so-called tight NLP (TNLP):
@@
We denote the feasible sets of the BNLP and TNLP by $\Omega_{\mathrm{TNLP}}$ and $\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}$, respectively. It can be seen that the following holds for $x^{*}\in\Omega$ Scheel2000 :
$\Omega_{\mathrm{TNLP}}=\bigcap_{(\mathcal{I}_{1},\mathcal{I}_{2})}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}\subset\Omega=\bigcup_{(\mathcal{I}_{1},\mathcal{I}_{2})}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}.$ (10)
From (10) for a feasible point $x^{*}\in\Omega$ the following can be concluded Scheel2000 . If $x^{*}$ is a local minimizer of the MPEC (1) then it is a local minimizer of the TNLP. The point $x^{*}$ is a local minimizer of the MPEC if and only if it is a local minimizer of every $\mathrm{BNLP}_{({\mathcal{I}}_{1},{\mathcal{I}}_{2})}$. The last assertion once again highlights the combinatorial nature of MPECs as – unless some strong assumptions specified below do not hold – an exponential number of branch NLPs must be checked to make conclusions about stationarity. The TNLP is used to define some MPEC-specific concepts, and we recall those relevant for this paper.
@@
with the MPEC Lagrange multipliers $\mu\in{\mathbb{R}}^{n_{c}}$, $\nu\in{\mathbb{R}}^{m}$ and $\xi\in{\mathbb{R}}^{m}$. By applying the KKT conditions to the TNLP (9) several stationarity concepts can be defined Scheel2000 , and they are listed in the following definition.
@@
  • Weak stationarity (W-stationarity) Scheel2000 : A point $x^{*}$ is called W-stationary if the corresponding TNLP (9) admits the satisfaction of the KKT conditions, i.e., there exist Lagrange multipliers $\mu^{*},\nu^{*}$ and $\xi^{*}$ such that:
@@
  • Mordukhovich stationarity (M-stationarity) Scheel2000 : A point $x^{*}$ is called M-stationary if it is weakly stationary and if either $\nu^{*}_{i}>0$ and $\xi^{*}_{i}>0$ or $\nu^{*}_{i}\xi^{*}_{i}=0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
@@
Note that in the simple case of ${\mathcal{I}}_{00}=\emptyset$, then all stationarity concepts are the same and they collapse to the concept of S-stationarity. From the definitions it follows that S-stationarity is the most restrictive since it has the tightest condition in the multipliers $\nu^{*}_{i},\xi^{*}_{i},i\in{\mathcal{I}}_{00}(x^{*})$. Weaker stationarity concepts, ordered by strength, are M, C, A, and W, with the stronger ones implying the weaker ones.
At this point, an interesting question is how are the multiplier-based stationarity concepts from Def. 2.3 related to the desired B-stationarity from Def. 2.1. The answer is given in the following theorem.
Theorem 2.2 (Theorem 4 in Scheel2000 )
If $x^{*}$ is a S-stationary point of the MPEC (1), then it is also B-stationary. If in addition, the MPEC-LICQ holds at $x^{*}$, then every B-stationary point is S-stationary.
As a consequence, if the restrictive MPEC-LICQ does not hold, there might exist B-stationary points that are not S-stationary. This assumption cannot be relaxed, as the next weaker condition, the MPEC-MFCQ is already not sufficient for S-stationarity, for a counterexample see (Scheel2000, , Example 3). Having a weaker termination criterion than B stationarity is arguably not satisfying as demonstrated by the next example due to Leyffer2007 .
Example 2.1
Consider the two-dimensional MPEC:
$\min_{x\in{\mathbb{R}}^{2}}\ (x_{1}-1)^{2}+x_{2}^{2}+x_{2}^{3}\quad\textnormal{s.t.}\quad 0\leq x_{1}\perp x_{2}\geq 0.$
The origin $x^{*}=0$ is an M-stationary point with the optimal multipliers $\nu=-2$, $\xi=0$. However, there exists a descent direction $d=(1,0)$ with $\nabla f(x^{*})^{\top}d=-2<0$. In conclusion, the origin is not B-stationary.
Even worse, despite MPEC-LICQ holding, both regularization  Ralph2004 ; Hoheisel2013 ; Kanzow2015 and some active set methods (cf. (Kirches2022, , Section 3)) can converge to points weaker than S-stationarity.
@@
In summary, without MPEC-LICQ, B-stationarity cannot be verified with multiplier-based stationarity concepts. We call all stationary points weaker than S-stationarity spurious stationary points. Moreover, even if MPEC-LICQ holds, most existing regularization methods and some active set methods may converge to points weaker than B-stationarity Hoheisel2013 . Given the drawbacks of multiplier-based stationarity concepts and the shortcomings of existing methods, it seems to be necessary to incorporate the LPEC (6) into the design of tailored algorithms to correctly certify B-stationary points. This is done in this paper.
2.2 Regularization and penalty-based methods
Regularization and penalty-based methods are closely related and can be used to find good starting points for our method. We briefly review some of the most effective ones. Although they can converge to spurious stationary points, they can be useful to quickly find a good initialization point for the algorithm introduced in this paper.
The family of regularization-based methods replaces the difficult complementarity constraints (2c) by a set of constraints parameterized by a parameter $\tau\geq 0$. The resulting NLPs are more regular, e.g., the MFCQ holds, and they can be solved efficiently with off-the-shelf NLP solvers. In practice, a (finite) sequence of such NLPs is solved while the parameter $\tau$ is driven towards zero. Surveys of such methods are available in Kanzow2015 ; Kim2020 ; Nurkanovic2024b .
One such method that performs well in practice Nurkanovic2024b is Scholtes’ global relaxation method Scholtes2001 , which solves a sequence of NLPs of the form of:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}$ $f(x)$ (Reg($\tau$))
s.t. $c(x)\geq 0,$
$x_{1},x_{2}\geq 0,$
$x_{1,i}x_{2,i}\leq\tau,$ $\forall i\in\{1,\ldots,m\}.$
We denote a stationary point of Reg($\tau$) by $x^{*}(\tau)$. Note that the feasible set of Reg($\tau$) contains the feasible set of (2). In practice, the first NLP $\mathrm{Reg}(\tau^{0})$ is solved usually with $\tau^{0}=1$ or $\tau^{0}=10$. Afterwards, the homotopy parameter $\tau^{k}$ is reduced via a geometric progression $\tau^{k+1}=\kappa\tau^{k}$ (e.g., $\kappa=0.1)$, and starting from $x^{*}(\tau^{k})$, the NLP Reg($\tau$) is resolved to obtain a new solution approximation $x^{*}(\tau^{k+1})$. If all problems Reg($\tau$) are feasible and MPEC-MFCQ holds, then the sequence of solutions $x^{*}(\tau^{k})$ converges to a C-stationary point as $\tau^{k}\downarrow 0$ Hoheisel2013 ; Scholtes2001 .
Another widely-used approach is the penalty reformulation Anitescu2005a ; Ralph2004 , which solve a sequence of NLPs:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}$ $f(x)+\frac{1}{\tau}\|\mathrm{diag}(x_{1})x_{2}\|$ (Pen($\tau$))
s.t. $c(x)\geq 0$
$x_{1},x_{2}\geq 0,$
We observe that the penalty functions $f(x)+\frac{1}{\tau}\|\mathrm{diag}(x_{1})x_{2}\|$ is smooth, as long as we ensure that all iterates remain nonnegative (can be readily enforced even in inexact arithmetics). In a practical implementation, one usually uses the $\ell_{1}$ or $\ell_{\infty}$ norm. For the former, the penalty function becomes $f(x)+\frac{1}{\tau}x_{1}^{\top}x_{2}$. The latter can be implemented with the help of a scalar slack variable Anitescu2005 .
Regularization and penalty-based methods are practical because they are easy to implement if we have a robust NLP solver. However, there are two major difficulties with these approaches:
  1. 1.
    Even for large values of the parameter $\tau$, the subproblems can become highly nonlinear or ill-conditioned, which may make them difficult to solve numerically.
  2. 2.
    Even if MPEC-LICQ holds, without stronger additional assumptions, relaxation-based methods may converge to spurious stationary points Hoheisel2013 ; Kanzow2015 ; Ralph2004 ; Scholtes2001 .
These difficulties are mitigated by the new algorithm presented in the next section. It solves mostly BNLPs, which are often easier than Reg($\tau$) and Pen($\tau$). Moreover, it always converges to B-stationary points, even if MPEC-LICQ does not hold.

### v2

@@ 2 Preliminaries
We start by defining some useful index sets for the complementarity constraints of the MPEC (1). Denote the feasible set of the MPEC (1) by $\Omega=\{x\in{\mathbb{R}}^{n}\mid c(x)\geq 0,0\leq x_{1}\perp x_{2}\geq 0\}$. The following index sets, which depend on a feasible point $x$ define a partition of $\{1,\ldots,m\}$:
${\mathcal{I}}_{0+}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}>0\},$
${\mathcal{I}}_{+0}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}>0,x_{2,i}=0\},$
${\mathcal{I}}_{00}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}=0\}.$
Most of the computational and theoretical difficulties arise if the so-called biactive set ${\mathcal{I}}_{00}(x)$ is nonempty. The active set of the standard inequality constraints (1b) is defined as ${\mathcal{A}}(x)=\{i\in\{1,\ldots,n_{c}\}\mid c_{i}(x)=0\}$}.
Next, we state first-order necessary optimality conditions. For this we need the notion of the tangent cone at $x\in\Omega$ to the set $\Omega$, which is defined as $\mathcal{T}_{\Omega}(x)=\{d\in{\mathbb{R}}^{n}\mid\exists\{x^{k}\}\subset\Omega,\{t^{k}\}\subset{\mathbb{R}}_{\geq 0}:\lim\limits_{k\to\infty}t^{k}=0,\lim\limits_{k\to\infty}x^{k}=x,\lim\limits_{k\to\infty}\frac{x^{k}-x}{t^{k}}=d\}$.
@@
((Luo1996, , Corollary 3.3.1)) Let $x^{*}\in\Omega$ be a local minimizer of (1), then it holds that
@@
or equivalently, $d=0$ is a global optimizer of the following optimization problem:
@@
Note that Theorem 2.1 is purely geometric and thus difficult to verify. In standard optimization theory, to obtain an algebraic characterization of a stationary point, the tangent cone ${\mathcal{T}}_{\Omega}(x)$ is replaced by the linearized feasible cone (Nocedal2006, , Def. 12.3). However, the latter is always a convex polyhedral cone, and if $\mathcal{I}_{00}(x)$ is nonempty the cone $\mathcal{T}_{\Omega}(x)$ is always nonconvex Luo1996 , making this transition impossible. A more suitable notion is the so-called MPEC-linearized feasible cone Flegel2005a ; Pang1999 ; Scheel2000 , which for a feasible point $x\in\Omega$ reads as
@@
Similarly to $x=(x_{0},x_{1},x_{2})$, the vector $d$ is partitioned as $d=(d_{0},d_{1},d_{2})\in\mathbb{R}^{p}\times\mathbb{R}^{m}\times\mathbb{R}^{m}$. Observe that at a feasible point $x\in\Omega$ all constraint residuals above are zero but they are kept for convenience. In contrast to the classical linearized feasible cone, here the combinatorial structure is kept for the degenerate index set ${\mathcal{I}}_{00}(x)$ and $\mathcal{F}_{\Omega}(x)$ is nonconvex, if $\mathcal{I}_{00}$ is nonempty. It is a union of a finite number of convex polyhedral cones.
In general, it holds that ${\mathcal{T}}_{\Omega}(x)\subseteq\mathcal{F}_{\Omega}(x)$ Flegel2005a . If the so-called MPEC-Abadie constraint qualification (MPEC-ACQ) is satisfied, i.e., ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$ Flegel2005a , we can use the $\mathcal{F}_{\Omega}(x)$ in Theorem 2.1. In particular, one can define a LPEC to characterize an algebraic B-stationary point (sometimes also called linearized B-stationarity Flegel2005a ), or B-stationary point for short.
Definition 2.1 (B-stationarity)
@@
s.t. $c(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,$ $\forall i\in{\mathcal{A}}(x^{*})$ (6b)
$x_{1,i}^{*}+d_{1,i}=0,\$$\forall i\in{\mathcal{I}}_{0+}(x^{*}),$ (6c)
$x_{2,i}^{*}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{*}),$ (6d)
$0\leq x_{1,i}^{*}+d_{1,i}\perp x_{2,i}^{*}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}).$ (6e)
To verify that $x^{*}$ is B-stationary, one has to verify that $d=0$ is a minimizer of an LPEC. Note that if $|{\mathcal{I}}_{00}(x^{*})|=0$, then (6) is an linear program (LP).
Let us define multiplier-based stationary concepts for MPECs and some auxiliary NLPs associated with the MPEC (1). Some of these problems will be solved by the algorithm introduced in this paper, while others serve to define additional concepts relevant to MPECs. To this end, we introduce further index sets.
Let $x^{*}\in\Omega$ and ${\mathcal{D}}_{1}(x^{*}),{\mathcal{D}}_{2}(x^{*})$ be a partition of the biactive set ${\mathcal{I}}_{00}(x^{*})$, with $\mathcal{D}_{1}(x^{*})\cup\mathcal{D}_{2}(x^{*})={\mathcal{I}}_{00}(x^{*})$ and $\mathcal{D}_{1}(x^{*})\cap\mathcal{D}_{2}(x^{*})=\emptyset$.
@@
Let us denote the set of all possible partitions by $\mathcal{P}(x^{*})=\{({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\}$. The number of possible partitions is $|\mathcal{P}(x^{*})|=2^{|{\mathcal{I}}_{00}(x^{*})|}$. For every partition, we can define a branch NLP, denoted by $\mathrm{BNLP}{({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))}$, as follows:
@@
$x_{1,i}\geq 0,\ x_{2,i}=0,\;$ $\forall i\in{\mathcal{I}}_{2}(x^{*}).$ (8d)
[…]
@@
[…]
$x_{1,i}\geq 0,\ x_{2,i}\geq 0,\;\forall i\in\mathcal{I}_{00}(x^{*}).$
[…]
It can be seen that the following holds for $x^{*}\in\Omega$ Scheel2000 :
$\Omega_{\mathrm{TNLP}}=\bigcap_{(\mathcal{I}_{1},\mathcal{I}_{2}){\in\mathcal{P}}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}\subset\Omega=\bigcup_{(\mathcal{I}_{1},\mathcal{I}_{2}){\in\mathcal{P}}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}.$ (10)
[…]
@@
with the MPEC Lagrange multipliers $\mu\in{\mathbb{R}}^{n_{c}}$, $\nu\in{\mathbb{R}}^{m}$ and $\xi\in{\mathbb{R}}^{m}$. By applying the KKT conditions to the TNLP (9), we obtain W-stationarity concept, and applying them to the RNLP yields S-stationary Scheel2000 . Further restrictions on the multipliers $\nu^{*}_{i},\xi^{*}_{i},\ i\in{\mathcal{I}}_{00}(x^{*})$, lead to M-, C-, and A-stationarity, which are not directly related to the KKT conditions of an auxiliary NLPs, see e.g.  Kim2020 ; Outrata1999 ; Scheel2000 . These concepts are summarized in the next definition.
@@
[…]
@@
  • Mordukhovich stationarity (M-stationarity) Outrata1999 : A point $x^{*}$ is called M-stationary if it is weakly stationary and if either $\nu^{*}_{i}>0$ and $\xi^{*}_{i}>0$ or $\nu^{*}_{i}\xi^{*}_{i}=0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
@@
We note that if ${\mathcal{I}}_{00}=\emptyset$, then all stationarity concepts are identical and collapse to the concept of S-stationarity. From the definitions it follows that S-stationarity is the most restrictive since it has the tightest condition in the multipliers $\nu^{*}_{i},\xi^{*}_{i},i\in{\mathcal{I}}_{00}(x^{*})$. Weaker stationarity concepts, ordered by strength, are M, C, A, and W, with the stronger ones implying the weaker ones. The relationship between multiplier-based stationarity (Def. 2.3) and B-stationarity (Def. 2.1) is established in the following theorem.
Theorem 2.2
((Scheel2000, , Theorem 4)) If $x^{*}$ is an S-stationary point of the MPEC (1), then it is also B-stationary. If, in addition, the MPEC-LICQ holds, then every B-stationary point is S-stationary.
@@
In summary, without the arguably restrictive MPEC-LICQ, B-stationarity cannot be verified using multiplier-based concepts. This assumption is essential, as the next weaker condition, MPEC-MFCQ, does not guarantee S-stationarity (cf. (Scheel2000, , Example 3)). Weaker termination criteria are unsatisfactory, since even M-stationarity allows trivial descent directions Leyffer2007 . Even under MPEC-LICQ, regularization Ralph2004 ; Hoheisel2013 ; Kanzow2015 and some active set methods (Kirches2022, , Section 3) may converge to points weaker than S-stationary. Therefore, we propose a method using the LPEC (6), to correctly certify B-stationarity.